# Source Code Vulnerability Detection: Combining Code Language Models and Code Property Graphs

### v1

@@ V-B Impact of the Tradeoff Parameter (RQ2)
The parameter $\lambda$ controls the trade-off between the training Vul-LMGNN and CodeBert. As $\lambda$ approaches 1, the model’s decisions rely more on the graph structure with the PL model embedding layer. Conversely, when $\lambda$ approaches 0, the model leans toward sequence-based decisions. Our experiments across different datasets reveal that the optimal value of $\lambda$ varies for different tasks, likely due to variations in vulnerability types and data distributions. For instance, on the partial Draper VDSIC dataset, increasing $\lambda$ does not significantly improve model performance. This phenomenon can be attributed to the strong performance of sequence-based methods on the former VDSIC dataset.
@@
Fig. 5 and 6 illustrate the evaluation matrix of Vul-LMGNN with varying $\lambda$ on the partial DiverseVul dataset. Accuracy improves consistently as $\lambda$ increases, reaching its peak at $\lambda$=0.8, slightly outperforming the GGNN or CodeBert alone (at $\lambda$ = 0 or 1). The achieved accuracy is 90.24%. During this process, precision exhibits fluctuations but overall shows an upward trend, reaching 46. 19% at $\lambda$ = 0.8, an improvement of 10. 71% over using the sequence model alone. However, recall follows a declining trend, reaching 36.45% at $\lambda$=0.8, indicating that transformer-based PL models exhibit higher recall in certain vulnerability detection scenarios.
V-C Evaluation of Fine-tuning Process (RQ3)
Pre-trained language models have demonstrated outstanding performance in various natural language processing tasks. Currently, there is a growing focus among researchers on employing pre-trained language models for code-related tasks, including code search, code completion, code summarization and so on [37, 13, 38, 39]. This has led to promising results in applications. This has prompted us to incorporate pre-trained models for programming languages in order to construct a novel vulnerability detection model.
We utilize the word embedding layer of pre-trained models as tokenization tools to generate node embeddings for graphs. These embeddings’ weights are further fine-tuned during training. In our experiments, we explore three settings. First, we initialize our embedding layer weights using a fine-tuned CodeBERT which perform fine-tuning on the target dataset [40]. Additionally, we compare this approach with two others: initializing embedding layer weights directly using pre-trained CodeBERT and GraphCodeBERT, respectively. The results are summarized in Table IV.
@@ V-C Evaluation of Fine-tuning Process (RQ3)
TABLE IV: Performance across various node embedding and initialization methods.
@@
Fine-tuned 84.38 87.37 80.64 83.87
Compared to directly using CodeBERT for node embedding weight initialization, GraphCodeBERT improves accuracy and recall by 0.3% and 0.89%, respectively. However, CodeBERT outperforms in precision and overall F1 score, achieving 87.75% and 83.61%. In contrast, using CodeBERT fine-tuned on the target dataset for embedding weight initialization yields the best overall performance, with the highest accuracy at 84.38% and an F1 score of 83.87%. Experimental results demonstrate that fine-tuning aids the pre-trained PL model in learning code embedding features specific to vulnerability distributions, further enhanced by GNNs for better detection performance.
V-D Different GNN Model Combination (RQ4)
To investigate the impact of combining different GNNs with pre-trained language models on vulnerability detection tasks, we compared three distinct GNNs: Graph Gated Neural Network (GGNN), Graph Convolutional Network (GCN) [41], and Graph Attention Network (GAT) [42], all integrated with CodeBERT. For a fair comparison, we followed the configuration from [37], maintaining a consistent two-layer GNN architecture and setting GAT’s number of heads to 8. Additionally, we employed fine-tuned CodeBERT with consistent model parameters. The specific experimental results are shown in Table V.
TABLE V: Performance across various GNN architectures.
@@ V-D Different GNN Model Combination (RQ4)
GGNN+CodeBERT 84.38 87.37 80.64 83.87
GCN+CodeBERT 83.08 86.90 77.90 82.15
GAT+CodeBERT 79.29 81.92 75.15 78.39
As shown in Table 5, our experiments were conducted on the partial Draper VDSIC dataset. The results indicate that GGNN exhibited the best overall performance, with an accuracy of 84.38% and an F1 score of 83.87%. Compared to GGNN, GCN experienced decreases in accuracy and precision by 1.3% and 0.47%, respectively, with the most significant decrease observed in recall at 2.74%. This may be attributed to GCN treating all neighboring nodes equally during convolution, thus failing to assign different weights based on node importance, leading to inaccurate identification of nodes related to code vulnerabilities. Additionally, GCN updates node features for the entire graph in a single computation, which poses challenges when dealing with complex code graph structures in inductive learning tasks related to code vulnerabilities.
@@
VI Conclusion
@@
In this paper, we propose a novel model, Vul-LMGNN, which integrates sequence and graph embedding techniques to detect vulnerabilities in function-level source code. Our approach leverages the code property graph representation of the source code as the primary input. Specifically, we utilize a pre-trained Program Language (PL) model to extract local semantic features from the code, which are then embedded as nodes in the graph using sequence-based embeddings. Subsequently, we employ a GGNN equipped with convolutional layers to effectively fuse heterogeneous information within the graph. Finally, our model jointly learns and predicts vulnerabilities by combining the PL model with the GGNN. To validate the effectiveness of Vul-LMGNN, we conducted extensive experiments on four real-world datasets, which demonstrated its superior performance. We systematically explored trade-off parameters, fine-tuning of the PL model, and variations of GNN architectures. Our findings further emphasize the positive contributions of each module to the overall model performance. As part of interesting future work, we intend to explore more effective fusion networks for learning code representations and facilitating multiclass detection.

### v2

@@ 5.2 Impact of the Tradeoff Parameter (RQ2)
The parameter $\lambda$ controls the trade-off between the training CodeLM and GNN. As $\lambda$ approaches 1, the model’s decisions rely more on the graph structure with the PL model embedding layer. Conversely, when $\lambda$ approaches 0, the model leans toward sequence-based decisions. Our experiments using Vul-LMGNNs (C-B) across various datasets indicated that the optimal value of $\lambda$ varies for different tasks, likely due to variations in vulnerability types and data distributions. For instance, on the partial Draper VDSIC dataset, increasing $\lambda$ does not significantly improve model performance. This phenomenon can be attributed to the strong performance of sequence-based methods on the former VDSIC dataset.
@@
Fig. 6 and 7 illustrated the evaluation matrix of Vul-LMGNNs with varying $\lambda$ on the partial DiverseVul dataset. Accuracy improves consistently as $\lambda$ increases, reaching its peak at $\lambda$=0.8, slightly outperforming the GGNN or CodeBert alone (at $\lambda$ = 0 or 1). The achieved accuracy is 90.24%. During this process, precision exhibits fluctuations but overall shows an upward trend, reaching 46. 19% at $\lambda$ = 0.8, an improvement of 10. 71% over using the sequence model alone. However, recall follows a declining trend, reaching 36.45% at $\lambda$=0.8, indicating that transformer-based PL models exhibit higher recall in certain vulnerability detection scenarios.
@@
Table 4: Performance across various node embedding and initialization methods.
@@
Fine-tuned(C-B) 84.38 87.37 80.64 83.87
Fine-tuned(GC-B) 84.51 86.36 81.97 84.11
5.3 Evaluation of Fine-tuning Process (RQ3)
Pre-trained LMs demonstrate outstanding performance in various natural language processing tasks. Currently, there is a growing focus among researchers on employing pre-trained LMs for code-related tasks, including code search, code completion, code summarization and so on [51, 17, 52, 53]. This has led to promising results in applications. This has prompted us to incorporate pre-trained models for programming languages in order to construct a novel vulnerability detection model.
We utilized the word embedding layer of pre-trained models as tokenization tools to generate node embeddings for graphs. These embeddings’ weights are further fine-tuned during training. In our experiments, we explored three settings. First, we initialized our embedding layer weights using a fine-tuned CodeBERT which perform fine-tuning on the target dataset [54]. Additionally, we compared this approach with two others: initializing embedding layer weights directly using pre-trained CodeBERT and GraphCodeBERT, respectively. The results are summarized in Table 4.
The experimental results revealed that fine-tuning Code LMs models improves performance over using the base models for node embedding initialization. the base model CodeBERT achieved an accuracy of 84.35%, with an F1 score of 83.61%. When employing GraphCodeBERT, the accuracy slightly decreased to 84.05%, while F1 score improved to 83.50%. In contrast, fine-tuned CodeBERT (C-B) model achieved an accuracy of 84.38%, an F1 score of 83.87%. Notably, the fine-tuned GraphCodeBERT (GC-B) model outperformed all others with an accuracy of 84.51%, and an F1 score of 84.11%. These results suggested that fine-tuning the models leads to a better overall balance in performance metrics.
5.4 Different GNN Model Combination (RQ4)
To investigate the impact of combining different GNNs with pre-trained LMs on vulnerability detection tasks, we compared distinct GNNs: Graph Gated Neural Network (GGNN), Graph Convolutional Network (GCN) [55], and Graph Attention Network (GAT) [56], and three variants of GraphSAGE (SAGE-mean, SAGE-GCN and SAGE-pool), all integrated with CodeBERT. For a fair comparison, we followed the configuration from [51], maintaining a consistent three-layer GNN architecture and setting GAT’s number of heads to 8. Additionally, we employed fine-tuned CodeBERT with consistent model parameters. The specific experimental results are shown in Table 5.
Table 5: Performance across various GNN architectures.
@@ 5.4 Different GNN Model Combination (RQ4)
GGNN+
CodeBERT 84.38 87.37 80.64 83.87
GCN+
CodeBERT 83.08 86.90 77.90 82.15
GAT+
CodeBERT 79.29 81.92 75.15 78.39
SAGE-mean+
CodeBERT 83.55 86.93 78.97 82.76
SAGE-GCN+
CodeBERT 84.18 87.05 80.31 83.54
SAGE-pool+
CodeBERT 81.90 85.82 76.43 80.85
As shown in Table 5, our experiments were conducted on the partial Draper VDSIC dataset. The results indicate that GGNN exhibit the best overall performance, with an accuracy of 84.38% and an F1 score of 83.87%. Compared to GGNN, GCN experiences decreases in accuracy and precision by 1.3% and 0.47%, respectively, with the most significant decrease observed in recall at 2.74%. This may be attributed to GCN treating all neighboring nodes equally during convolution, thus failing to assign different weights based on node importance, leading to inaccurate identification of nodes related to code vulnerabilities. Additionally, GCN updates node features for the entire graph in a single computation, which poses challenges when dealing with complex code graph structures in inductive learning tasks related to code vulnerabilities.
@@
Table 6: Ablation study for online knowledge distillation; "self" denotes the absence of online knowledge distillation.
KD Ablation ACC(%) P(%) R(%) F1(%)
Self 83.50 86.68 79.17 82.75
2 Student 84.38 87.37 80.64 83.87
3 Student 84.18 84.77 83.33 84.04
4 Student 83.90 87.93 78.59 83.00
5 Student 84.15 88.93 78.01 83.11
6 Student 83.61 86.22 80.01 83.00
5.5 Evaluation of Online KD for GGNNs (RQ5)
Figure 4 reveals the flow of structural information during the iterative training process of two GNNs in the online knowledge distillation approach. It can be observed that each layer of one student model aligns with the next layer of the other student model, with the final layer particularly matching the first layer. In this way, the two student models exchange structural information with each other, and propagate it across different layer depths. In this work, we evaluated two aspects of this approach: 1) the performance gain achieved and 2) how the performance changes as the number of student models is varied.
When transitioning from the "Self" baseline, which denotes the absence of online knowledge distillation, to using two student models, there is a significant enhancement across all measured metrics. This initial improvement underscores the benefit of the distillation process. Further analysis shows that the performance continues to improve with the addition of a third student model. However, this trend does not persist uniformly as the number of student models increases beyond three. For instance, while the metrics improve when moving from two to three student models, the subsequent addition of a fourth student model does not yield substantial gains and is followed by a slight decline in Accuracy and Precision. This indicates that, after reaching a certain threshold, the benefits of adding more student models begin to diminish.
5.6 Compared Online KD with MixHop
In this paper, we leverage Online KD to address the limitations of single-layer information propagation in GNNs. This approach enables our method to effectively capture the intricate logic within abstract code graphs. While acknowledging the existence of algorithms like MixHop[57] in the GNN literature, which aim to capture long-range dependencies in graph neural networks, our experiment seeks to analyze and compare our KD-based method with MixHop.
We begin by examining the fundamental differences between these two approaches. MixHop is designed to overcome GNNs’ limitations in handling long-distance relationships between nodes by mixing node features across multiple hop counts (i.e., neighborhood layers). In contrast, our Online Knowledge Distillation method focuses on transferring knowledge between teacher and student models. The key distinction is that online KD facilitates knowledge transfer between teacher and student models, whereas MixHop integrates node features across multiple hop counts. The implementation results are presented in the Table7, providing a comparative analysis of these two approaches in addressing the challenges of information propagation and logic capture in GNNs for abstract code graphs.
Based on the experimental results, the Knowledge Distillation (KD) based method demonstrates superior performance compared to the MixHop-based approach on the Partial Draper VDSIC Dataset. The KD-based method achieves higher accuracy (84.38% vs 83.68%) and significantly better recall (80.64% vs 78.47%), indicating its enhanced ability to identify relevant instances. The KD-based approach’s higher F1 score (83.87% vs 82.78%) underscores its better overall balance between precision and recall. These results suggest that the KD-based method is more effective at capturing and utilizing relevant information from the dataset, making it the preferable choice.
Table 7: Comparison of Experimental Results for MixHop-based Methods and KD-based Methods on the Partial Draper VDSIC Dataset
Methods ACC(%) P(%) R(%) F1(%)
MixHop-based 83.68 87.60 78.47 82.78
KD-based 84.38 87.37 80.64 83.87
6 Conclusion
@@
In this paper, we propose a novel model, Vul-LMGNNs, which integrates sequence and graph embedding techniques to detect vulnerabilities in function-level source code. Our approach leverages the code property graph representation of the source code as the primary input. Specifically, we utilize a pre-trained Program Language (PL) model to extract local semantic features from the code, which are then embedded as nodes in the graph using sequence-based embeddings. Subsequently, we employ a GGNN equipped with convolutional layers to effectively fuse heterogeneous information within the graph. Finally, our model jointly learns and predicts vulnerabilities by combining the PL model with the GGNN. To validate the effectiveness of Vul-LMGNNs, we conducted extensive experiments on four real-world datasets, which demonstrated its superior performance. We systematically explored trade-off parameters, fine-tuning of the PL model, and variations of GNN architectures. Our findings further emphasize the positive contributions of each module to the overall model performance. As part of interesting future work, we intend to explore more effective fusion networks for learning code representations and facilitating multiclass detection.